OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle with legs of 9cm and 12 cm

OpenStudy (anonymous):

Use a^2 + b^2= c^2 So a= 9 and b= 12 To find c, square root the sum of the squares of 9 and 12. ;)

OpenStudy (gorv):

use pythagorse therom (leg1)^2+(leg2)^2=hypotenuse^2

OpenStudy (hba):